2. Water of m = 5 kg m , P = 1 MPa, H = 10,280 kJ. Determine temperature, total volume, total

internal energy and total entropy.


1
Expert's answer
2021-05-07T13:24:02-0400

specific enthalpy h = H/m = 10,280kJ/5kg = 2056 kJ/kg

internal energy u = h - pv = 2056kJ/kg - (10^6 N/m^2)*(specific volume (m3/kg))